Magnetic field at the centre of various regular polygons which are made by a constant length l current carrying wire as regular hexagon is???(assume side length a and direction of current anticlockwise)

Grade:12

1 Answers

Khimraj
5 years ago
Magnetic field at centre of n sided regular polygen is equal to n times of magnetic field due to one side.
B = n(uoI/4\pir)(sin\Theta 1 + sin\Theta 2)
where r = acos\Theta and \Theta1 = \Theta2
So B  = (nuoI/2\pia)(tan\Theta)
\Theta = \pi/n
where n is no side in regular polygen
Then
B = (nuoI/2\pia)(tan(\pi/n))
Hope it clears. If you like anser then please approve it.